### v2

# Broken intrinsic symmetry induced magnon-magnon coupling in synthetic ferrimagnets

Abstract
Synthetic antiferromagnets offer rich magnon energy spectra in which optical and acoustic magnon branches can hybridize. Here, we demonstrate a broken intrinsic symmetry induced coupling of acoustic and optical magnons in a synthetic ferrimagnet consisting of two dissimilar antiferromagnetically interacting ferromagnetic metals. Two distinct magnon modes hybridize at degeneracy points, as indicated by an avoided level-crossing. The avoided level-crossing gap depends on the interlayer exchange interaction between the magnetic layers, which can be controlled by adjusting the non-magnetic interlayer thickness. A large avoided level crossing gap of 3.9 GHz is revealed, exceeding the coupling strength that is typically found in other magnonic hybrid systems based on a coupling of magnons with photons or magnons with phonons.
I Introduction
Magnons are the quanta of spin waves representing the fundamental excitation of the magnetic order in a magnetic system. Magnons play a crucial role in our understanding of magnetism. They have been the subject of extensive study for their properties and applications in various technologies, particularly in the development of novel magnonic devices Chen et al. (2021); Haldar et al. (2016); Urazhdin et al. (2014); Khitun et al. (2010); Mahmoud et al. (2020); Chumak et al. (2022); Wang et al. (2024a). Antiferromagnetic magnons are investigated for their potential use in ultrafast applications including next-generation memory devices and information processing Xiong et al. (2022); Kosub et al. (2017); Olejník et al. (2017, 2018); Jungfleisch et al. (2017), terahertz spectroscopy and imaging Lee et al. (2021); Rongione et al. (2023); Mashkovich et al. (2019); Bowlan et al. (2018); Wu et al. (2021), and quantum computing Meier et al. (2003). Giant magnetoresistance (GMR) was discovered in 1988 in antiferromagnetically coupled layered ferromagnetic (FM) structures Baibich et al. (1988); Binasch et al. (1989). This marks one of the initial applications of synthetic antiferromagnetic (sAF) heterostructures comprising two ferromagnetic metal layers coupled antiferromagnetically. This idea expanded to various other devices such as magnetic tunnel junctions and spin-torque nano-oscillators Grollier et al. (2016); Kiselev et al. (2003); Luo et al. (2023). sAFs offer several engineering advantages compared to bulk antiferromagnets. Unlike bulk antiferromagnetic materials, sAFs can be tailored with specific layer thicknesses and compositions, which allows for the optimization of magnetic properties such as exchange bias, coercivity, and anisotropy. These engineering opportunities make sAFs ideal model systems for studying antiferromagnetic magnons.
Coupling between magnon modes in sAFs can lead to the exchange of energy and momentum between different magnon modes. By controlling the coupling strength, the propagation and transmission of spin waves can be manipulated. This can improve performance in magnonic devices such as magnon transistor Gong et al. (2018), and magnon logic elements Chumak (2019). The most dominant mechanisms responsible for magnon-magnon coupling Zhang et al. (2025) are the exchange interaction (short range) MacNeill et al. (2019) and the dipolar interaction (long range) Shiota et al. (2020). In sAFs comprising two sublattices with antiferromagnetic exchange interaction, two distinct normal modes can emerge based on the relative motion of the spins precessing in the sublattices; in-phase motion (acoustic mode) or out-of-phase motion (optical mode). Generally, these modes are protected by symmetry, preventing them from coupling Neumann and Wigner (1929); Patchett et al. (2022). In a symmetric synthetic antiferromagnet, where the two ferromagnetic layers are magnetically equivalent, a $180^{\circ}$ rotation about the field direction (when the field is aligned with a high-symmetry axis) is a valid symmetry operation. Under this operation, the acoustic mode is odd ($A\rightarrow-A$) and the optical mode is even ($A\rightarrow A$). Any coupling term must be invariant under the symmetry operation; however, the product of an even and an odd mode changes sign and therefore the coupling vanishes, allowing the modes to cross without hybridization. Strategies to facilitate magnon-magnon coupling in sAFs have been studied recently Sud et al. (2020); Dai and Ma (2021); Hayashi et al. (2023); Shiota et al. (2020); Sklenar and Zhang (2021); Chen et al. (2020).
Furthermore, a recent theoretical work suggested the possibility of magnon-magnon coupling for an in-plane geometry in an intrinsic symmetry-breaking mechanism using a synthetic ferrimagnetic structure Li et al. (2021). Ferrimagnets are partially compensated magnetic structures in which the sublattices with unequal magnetic moments are arranged in opposite directions, which results in non-zero spontaneous magnetization. In a synthetic structure, this can be achieved in different ways, such as by using FM layers with different saturation magnetization, $M_{\mathrm{s}}$, by using different thicknesses of the FM layers, $d$, by using an unequal number of layers, or any combination of them, all connected by negative Ruderman–Kittel–Kasuya–Yosida (RKKY) 111also known as bilinear or quadratic interaction interaction between layers Lepadatu et al. (2017); Xie et al. (2023); Jenkins et al. (2014); Franco and Landeros (2018); Sud et al. (2023); Wang et al. (2024b); Subedi et al. (2024). Synthetic ferrimagnets can facilitate intrinsic symmetry breaking in the presence of RKKY and biquadratic interactions, which leads to the coupling of acoustic and optical modes Li et al. (2021).
Figure 1: (a) Schematic illustration of the synthetic ferrimagnet comprising of a CoFe/Ru/NiFe heterostructure. The thickness of Ru is varied from 0 nm to 11 nm to tune the coupling strength between CoFe and NiFe. (b) Illustration of the spin-torque ferromagnetic resonance (STFMR) setup, including an inset that shows the relevant torques on the magnetization $\boldsymbol{M}$: $\boldsymbol{\tau}_{||}$ - anti-damping-like torque, $\boldsymbol{\tau}_{\alpha}$ - damping-like torque, $\boldsymbol{\tau}_{\perp}$ - field-like torque. The biasing field $\boldsymbol{H}_{0}$ is applied at an angle of 45 degrees with respect to the microwave field $\boldsymbol{H}_{\mathrm{rf}}$. (c) Schematic diagram of equilibrium angles of the ferromagnetic layers relative to the external field direction. (d) Vibrating sample magnetometry (VSM) measurement for a sample with Ru thickness of $0.8$ nm. Corresponding macrospin model calculations and micromagnetic simulations (MuMax3) are overlaid. Both the macrospin model and simulations show good agreement with experiment. The arrows indicate the relative alignment of the two exchange-coupled ferromagnetic layers. (e) MuMax3 simulations of the equilibrium angles of the ferromagnetic layers as a function of the applied field during field ramp-down, referenced to the external field direction defined in (c). Note that -180° and 180° correspond to the same magnetization direction. In some regions, the blue data points (magnetization direction of CoFe) are obscured by the overlapping orange data points (magnetization direction of NiFe).
Here, we report the experimental observation of such magnon-magnon coupling induced by broken intrinsic symmetry in synthetic ferrimagnetic structures. We use two antiferromagnetically interacting magnetic layers of similar thickness, $d$, but different saturation magnetization, $M_{\mathrm{s}}$ (CoFe and NiFe) separated by a thin spacer layer (Ru). Magnetometry using vibrating sample magnetometry (VSM) reveals the exchange interaction between the ferromagnetic layers. Spin dynamics spectrum reveals that the asymmetry introduced by broken intrinsic symmetry facilitates the coupling between the acoustic and optical magnon modes. Furthermore, our studies show that the coupling strength between the modes can be tuned by controlling the exchange interaction between the ferromagnetic layers. Our experimental results are systematically compared with macrospin modeling and micromagnetic simulations using MuMax3 Vansteenkiste et al. (2014). Our study aims to leverage the magnon-magnon coupling within an in-plane geometry to develop innovative magnonic device applications.
II Sample Fabrication and Experimental Details
Figure 1(a) illustrates a schematic of the synthetic ferrimagnetic sample stack. The samples comprise a trilayer of $\mathrm{Co_{90}}\mathrm{Fe_{10}(10\penalty 10000\ nm)}$$\mathrm{Ru(}d\penalty 10000\ \mathrm{nm)}$$\mathrm{Ni_{80}}\mathrm{Fe_{20}(10\penalty 10000\ nm)}$ with varying thickness $d$ of the non-magnetic spacer layer Ru. The structure was grown on thermally oxidized Si substrates at room temperature under an Ar pressure of $\mathrm{4.5\times{10^{-3}}}$ Torr in a magnetron sputtering system with a base pressure of $\mathrm{1.4\times{10^{-7}}}$ Torr. The Ru spacer layer was grown in a wedge shape, allowing for a continuous thickness variation from $d=0$ to 1 nm at an wedge angle of about $\mathrm{1\penalty 10000\ }\mathrm{\mu}\mathrm{deg}$, by off-center deposition to tune the interlayer exchange interaction. All samples were capped with a 5 nm $\mathrm{SiO}_{\mathrm{2}}$ layer to protect the sample from oxidation. The sample was then cut into 11 rectangular samples where the thickness of the spacer layer can be considered approximately constant over the width of each sample. The anisotropic easy axes in the FM layers were introduced along the long sides of the samples. VSM measurements were performed to characterize the static magnetic properties of the individual pieces of the sample. Figure 1(d) shows a typical VSM measurement of a synthetic ferrimagnet with negative RKKY interaction between the ferromagnetic layers, which we will discuss below.
To probe the spin dynamics of the ferrimagnet, we employ spin-torque ferromagnetic resonance (STFMR) spectroscopy. The STFMR signal is the spin rectification of the magnetoresistance in the device Gui et al. (2007), which provides high sensitivity measurements. In STFMR measurements a spin-polarized current in addition to an RF magnetic field exerts a torque on the magnetization, causing it to precess when the resonance condition is realized. This precession induces a time-varying magnetization, consequently giving rise to an AC resistance as a result of the anisotropic magnetoresistance or spin Hall magnetoresistance Tulapurkar et al. (2005); Sankey et al. (2006). This, combined with the AC current passed through the sample at the same frequency, generates a rectified DC voltage, which is measured using a lock-in amplifier Liu et al. (2011). For STFMR measurements, the sputtered thin films were etched to produce STFMR devices following the following steps: first, a negative-tone photoresist was used to cover 80 $\mathrm{\mu}$m $\times$ 130 $\mathrm{\mu}$m rectangular sections of the film, followed by etching using Ar ion milling of the surrounding film. The shorted coplanar waveguides were then patterned on top of the rectangular sections using optical lithography and positive photoresist. Finally, 5 nm of Ti and 20 nm of Au were e-beam evaporated and lifted off to create the waveguides on the devices as shown in Fig. 1(b).
Figure 2: (a,b) Experimentally measured magnon spectra for Ru thicknesses of 0.8 nm and 0.4 nm, with corresponding macrospin model calculations overlaid. The color scale represents the lock-in amplifier output. (c,d) Corresponding micromagnetic (MuMax3) simulations of the STFMR spectra, with macrospin model results shown for comparison. Here, the color scale represents the FFT amplitude.
A bias tee was employed to simultaneously apply a microwave current and measure the rectified DC voltage using a lock-in amplifier Hossain et al. (2021). A microwave signal with an output power of +21 dBm was generated by an Agilent E8257D signal generator and amplitude-modulated at a frequency of 666.7 Hz. The external magnetic field, $\boldsymbol{H}_{\mathrm{0}}$, was applied in the plane of the sample at an angle of $45^{\circ}$ relative to the device axis [see Fig. 1(b)] to maximize the signal amplitude Harder et al. (2011); Liu et al. (2011). For each applied field, the magnetic moments of the two ferromagnetic layers (NiFe and CoFe) reach equilibrium orientations characterized by angles $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$, respectively, with respect to the external field direction $\boldsymbol{H}_{0}$ [see Fig. 1(c)]. These equilibrium angles were determined from micromagnetic simulations [see Fig. 1(e)]. The microwave frequency was swept from 1 GHz to 18 GHz in 86 steps, while the external magnetic field was varied from +200 mT to 0 mT for each frequency step.
Figure 3: Micromagnetic simulations obtained with MuMax3. (a) Simulated spectrum for a Ru thickness of 0.8 nm. (b–f) Layer-resolved simulations showing the amplitude (top) and phase (bottom) profiles of the magnon modes, extracted from FFTs of the magnetization dynamics at 4, 12, 37, 56, and 100 mT. The two vertical lines in each panel indicate the resonance frequencies at that field. These lines intersect the phase spectra of CoFe (black) and NiFe (red) at the lower-frequency (solid) and higher-frequency (dashed) branches, highlighted by the corresponding horizontal markers in the phase panels. (g) Spatially resolved simulation at 12 mT with X and Y indicating the lateral dimensions of the studied devices. The color wheel represents the magnetization dynamics phase.
III Results and Discussion
As is shown in Fig. 1(d), magnetometry measurements show a typical hysteresis loop of a synthetic antiferromagnet. Magnetometry measurements show secondary loops in the hysteresis for Ru thicknesses above $0.2$ nm [see Fig. S2 in the supplementary materials (SM)]. From 0.20.50.2-0.5 nm, two secondary loops appear near the knee (the field at which the magnetization changes from a saturated to an unsaturated state), while above 0.5 nm, a loop forms at the center [Fig. 1(d)]. The ‘knee loops’ suggest the presence of strong biquadratic exchange interaction between FM films for Ru thicknesses of 0.20.50.2-0.5 nm Demokritov (1998) (see SM). The magnetometry results can be explained by carefully analyzing the expression of the volume energy density of the structure Inoue (1994):
$\mathrm{\mathscr{E}}=-\boldsymbol{\mathrm{M}}\cdot\boldsymbol{\mathrm{H_{0}}}+J_{\mathrm{q}}\boldsymbol{m}_{1}\cdot\boldsymbol{m}_{2}+J_{\mathrm{{bq}}}(\boldsymbol{m}_{1}\cdot\boldsymbol{m}_{2})^{2}.$ (1)
Here, $\boldsymbol{M}$ is the magnetization, $\boldsymbol{H}_{\mathrm{0}}$ is the external magnetic field, and $\boldsymbol{m}_{\mathrm{i}}$ are the magnetization unit vectors of the $i$-th layer ($i=1,2$), respectively. $J_{\mathrm{q}}$ and $J_{\mathrm{{bq}}}$ are the phenomenological energy densities for quadratic and biquadratic exchange interaction. RKKY interaction is negative (i.e., favoring antiferromagnetic alignment of the magnetic layers) when $J_{\mathrm{q}}$ is positive. At a large field, the magnetostatic energy is large enough to saturate both layers in the direction of the field [configuration labeled ‘1’ in Fig. 1(d)]. As the field is reduced, a field is reached [beginning of configuration ‘2’ in Fig. 1(d)] where the exchange energy becomes significant and attempts to align the FM layers antiferromagnetically. As a result, the two layers reach an equilibrium at a tilted angle with respect to each other away from the saturated state, reducing the net magnetization. The critical field at which the magnetization starts to reduce (i.e., the knee field in the broader hysteresis loop) can be calculated using a macrospin model as in Eq. (1) (see SM for the derivation):
$\mathrm{H_{\mathrm{{0,ff}}}=(J_{\mathrm{q}}+2J_{\mathrm{{bq}}})\left(\frac{1}{M_{\mathrm{{S1}}}}+\frac{1}{M_{\mathrm{{S2}}}}\right)},$ (2a)
where $H_{\mathrm{{0,ff}}}$ is the magnitude of the magnetic field over which the FM layers align in a forced ferromagnetic state [configuration ‘1’ in Fig. 1(d)], and $M_{\mathrm{Si}}$ are the saturation magnetizations of the $i$-th layer ($i=1,2$), respectively.
If the biquadratic exchange interaction is negligible, magnetization changes linearly with the external field in the unsaturated configuration [configuration ‘2’ in Fig. 1(d)] (see SM for derivation). However, strong biquadratic interaction causes a non-linear relation, creating secondary ‘knee loops’ near the knee field in the hysteresis (see SM).
As the external field decreases further, the quadratic exchange energy dominates, forming a plateau [configuration ‘3’ in Fig. 1(d)] with antiparallel layer alignment. The finite magnetization at this plateau results from different saturation magnetizations between the ferromagnetic layers Waring et al. (2020), creating the ‘central loop’ in the magnetometry data, as seen in Fig. 1(d). The critical field at which the layers align antiferromagnetically (i.e., the plateau region in the central part of the hysteresis loop) is given by Eq. (2b):
$H_{\mathrm{0,afm}}=(J_{\mathrm{q}}-2J_{\mathrm{bq}})\left(\frac{1}{M_{\mathrm{S2}}}-\frac{1}{M_{\mathrm{S1}}}\right),$ (2b)
where, when the external magnetic field lies between $H_{\mathrm{0,afm}}$ and $H_{\mathrm{0,afm}}$, the ferromagnetic layers align antiferromagnetically [configuration ‘3’ in Fig. 1(d)], and $M_{\mathrm{S1}}>M_{\mathrm{S2}}$. Equation (2b) explains the absence of a central loop in magnetometry when Ru thickness is less than $0.6$ nm, given $J_{\mathrm{q}}<2J_{\mathrm{bq}}$ makes $H_{\mathrm{0,afm}}$ negative, preventing a fully antiferromagnetic state. Eqs. (2a) and (2b) provide a means to extract $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ from magnetometry measurements when $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, allowing a fully antiferromagnetic state:
$\mathrm{J_{q}=\left(\frac{H_{0,ff}}{M_{S1}+M_{S2}}+\frac{H_{0,afm}}{M_{S1}-M_{S2}}\right)\frac{M_{S1}M_{S2}}{2}},$ (3a)
$\mathrm{J_{bq}=\left(\frac{H_{0,ff}}{M_{S1}+M_{S2}}-\frac{H_{0,afm}}{M_{S1}-M_{S2}}\right)\frac{M_{S1}M_{S2}}{4}}.$ (3b)
The relationship between the saturation magnetizations of the ferromagnetic layers can be established using the normalized magnetization of the forced ferromagnetic ($M_{\mathrm{ff}}$) and fully antiferromagnetic ($M_{\mathrm{afm}}$) states (the magnetizations at configuration ‘1’ and configuration ‘3’ in Fig. 1(d) respectively). The ratio of the saturation magnetizations can be expressed as follows:
$M_{\mathrm{r}}=\mathrm{\frac{M_{S2}}{M_{S1}}=\frac{M_{ff}-M_{afm}}{M_{ff}+M_{afm}}}.$ (4)
For a Ru thickness of $0.8\,\mathrm{nm}$, we extracted the critical fields from the VSM results [Fig. 1(d)] as $H_{0,\mathrm{ff}}=55\pm 1.5\,\mathrm{mT}$ and $H_{0,\mathrm{afm}}=13\pm 1\,\mathrm{mT}$, with corresponding normalized magnetizations $M_{\mathrm{ff}}=1.0$ and $M_{\mathrm{afm}}=0.302\pm 0.028$. From these values, the ratio of the saturation magnetization is calculated as $M_{\mathrm{r}}=0.536\pm 0.033$.
Figures 2(a,b) show the STFMR spectra for samples with Ru thicknesses of 0.8 nm and 0.4 nm, both exhibiting antiferromagnetic exchange interaction as confirmed by magnetometry (see Fig. S2). We notice two magnon modes in the spectra of all of our samples with negative RKKY interaction (as found in magnetometry measurements; see also SM). These two modes correspond to the optical and acoustic modes (more details on the phase of the modes can be found below). Typically, these two modes are protected by symmetry, which means that they cross at the degeneracy point. As is evident from Fig. 2, that is not what we observe in the experiments: a clear avoided-level crossing is observed at the degeneracy points, indicating strong coupling between the optical and acoustic modes. The coupling gap is determined by fitting the resonance peaks in the STFMR spectra with standard Lorentzian functions (see SM for details) and identifying the minimum frequency separation between the resonances at a fixed magnetic field. The largest gap of 3.9 GHz is observed for the sample with a Ru thickness of 0.4 nm. The enhanced signal amplitude at 3.8 GHz is a measurement artifact and does not carry physical significance.
We model the magnetization dynamics by treating the two ferromagnetic layers as coupled macrospins, whose time evolution is governed by the Landau–Lifshitz–Gilbert (LLG) equation Landau and Lifshitz (1935); Gilbert (2004); Stancil and Prabhakar (2008):
$\frac{d\mathbf{m}_{\mathrm{1}}}{dt}$ $=-\mu_{0}\gamma\mathbf{m}_{\mathrm{1}}\times\Big[H_{0}\hat{\mathbf{y}}-H_{\mathrm{E1}}\mathbf{m}_{\mathrm{2}}$
$\qquad-H_{\mathrm{B1}}(\mathbf{m}_{\mathrm{1}}\cdot\mathbf{m}_{\mathrm{2}})\mathbf{m}_{\mathrm{2}}-M_{\mathrm{S1}}(\mathbf{m}_{\mathrm{1}}\cdot\hat{\mathbf{z}})\hat{\mathbf{z}}\Big$]. (5a)
$\frac{d\mathbf{m}_{\mathrm{2}}}{dt}$ $=-\mu_{0}\gamma\mathbf{m}_{\mathrm{2}}\times\Big[H_{0}\hat{\mathbf{y}}-H_{\mathrm{E2}}\mathbf{m}_{\mathrm{1}}$
$\qquad-H_{\mathrm{B2}}(\mathbf{m}_{\mathrm{1}}\cdot\mathbf{m}_{\mathrm{2}})\mathbf{m}_{\mathrm{1}}-M_{\mathrm{S2}}(\mathbf{m}_{\mathrm{2}}\cdot\hat{\mathbf{z}})\hat{\mathbf{z}}\Big$]. (5b)
Here, $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ denote the macrospin unit vectors of the ferromagnetic layers, with their dynamics described by
$\frac{d\mathbf{m}_{\mathrm{i}}}{dt}=i\omega\,\delta\mathbf{m}_{\mathrm{i}},$
where $\delta\mathbf{m}_{\mathrm{i}}$ is the unit vector of the dynamic deviation of the magnetization unit vector. The effective fields $H_{\mathrm{E\,i}}$ and $H_{\mathrm{B\,i}}$ due to quadratic and biquadratic exchange are defined as
$H_{\mathrm{E\,i}}=\frac{J_{\mathrm{q}}}{M_{\mathrm{S\,i}}},\qquad H_{\mathrm{B\,i}}=\frac{2J_{\mathrm{bq}}}{M_{\mathrm{S\,i}}}.$
Finally, the torque terms proportional to the saturation magnetizations $M_{\mathrm{S1}}$ and $M_{\mathrm{S2}}$ in Eqs. 5a and 5b parameterize the shape-dependent demagnetizing fields of the NiFe and CoFe layers, respectively.
After some algebra (see SM), the nontrivial solutions are obtained by requiring that the determinant of the coefficient matrix vanishes:
$\begin{vmatrix}i\omega&\mu_{0}\gamma(M_{1}+H_{\mathrm{1,eq}})&0&\frac{\mu_{0}\gamma}{M_{1}}(J_{\mathrm{q}}+2J_{\mathrm{bq}}\cos{\alpha})\\ -\mu_{0}\gamma\left(\frac{2J_{\mathrm{bq}}}{M_{1}}\sin^{2}{\alpha}+H_{\mathrm{1,eq}}\right)&i\omega&\frac{\mu_{0}\gamma}{M_{1}}\left[2J_{\mathrm{bq}}(\sin^{2}{\alpha}-\cos^{2}{\alpha})-J_{\mathrm{q}}\cos{\alpha}\right]&0\\ 0&\frac{\mu_{0}\gamma}{M_{2}}(J_{\mathrm{q}}+2J_{\mathrm{bq}}\cos{\alpha})&i\omega&\mu_{0}\gamma(M_{2}+H_{\mathrm{2,eq}})\\ \frac{\mu_{0}\gamma}{M_{2}}\left[2J_{\mathrm{bq}}(\sin^{2}{\alpha}-\cos^{2}{\alpha})-J_{\mathrm{q}}\cos{\alpha}\right]&0&-\mu_{0}\gamma\left(\frac{2J_{\mathrm{bq}}}{M_{2}}\sin^{2}{\alpha}+H_{\mathrm{2,eq}}\right)&i\omega\\ \end{vmatrix}=0,$
where $H_{\mathrm{i,eq}}$ denotes the effective magnetic field experienced by layer $i$, and $\alpha=\phi_{\mathrm{a}}+\phi_{\mathrm{b}}$ is the equilibrium interlayer angle.
Numerical solutions of the determinant yield the magnon branches for a given set of parameters, and the experimental spectra can be fitted within this framework to extract the corresponding material parameters. Fitting the STFMR spectra with the macrospin model is complicated by the hidden variables $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$, which describe the equilibrium magnetization orientations of the two ferromagnetic layers and cannot be measured directly. Experimentally, we obtain the resonance frequency as a function of the external magnetic field, while the angles $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$ must be determined self-consistently from the model. The fitting algorithm and corresponding Python code are provided in the SM.
For a Ru thickness of $0.8\,\mathrm{nm}$, the fitted parameters are $M^{\mathrm{fit}}_{\mathrm{S1}}=1.82\times 10^{6}\penalty 10000\ \mathrm{A/m}$, $M^{\mathrm{fit}}_{\mathrm{r}}=0.527$, $J^{\mathrm{fit}}_{\mathrm{q}}=29.8\penalty 10000\ \mathrm{kJ/m^{3}}$, and $J^{\mathrm{fit}}_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$. The fitted macrospin magnon spectrum is overlaid with the STFMR data in Fig. 2(a). An analogous procedure was followed for the sample with a Ru thickness of $0.4\penalty 10000\ \mathrm{nm}$, shown in Fig. 2(b), demonstrating good agreement in both cases. Using the material parameters obtained from the dynamic macrospin model, we compare the calculated hysteresis loop with the experimental VSM data shown in Fig. 1(d) and find good overall agreement. Finally, these parameters also yield the knee fields from Eqs. (2a) and (2b), resulting in $H^{\mathrm{fit}}_{0,\mathrm{ff}}=54.4\penalty 10000\ \mathrm{mT}$ and $H^{\mathrm{fit}}_{0,\mathrm{afm}}=12.6\penalty 10000\ \mathrm{mT}$, which are consistent with the values extracted from the VSM measurements ($H_{0,\mathrm{ff}}=55\pm 1.5\,\mathrm{mT}$ and $H_{0,\mathrm{afm}}=13\pm 1\,\mathrm{mT}$) within experimental uncertainty.
To gain a better understanding of the experimental results, we performed micromagnetic modeling using the graphics processor unit (GPU)-accelerated program MuMax3 Vansteenkiste et al. (2014); Exl et al. (2014). Our structure was modeled as $128\times 64\times 3$ cells with an individual cell size of $4.0$ nm × $\times\penalty 10000\ 4.0$ nm × $\times\penalty 10000\ 4.0$ nm, with the middle layer representing the spacer. Each layer was defined into a region to set material parameters separately. The values for the exchange stiffness of CoFe and NiFe used in the simulation were $3.0\times 10^{-11}$ J/m and $1.3\times 10^{-11}$ J/m, respectively Parreiras and Martins (2015). This corresponds to exchange lengths for NiFe and CoFe of 5.29 nm and 4.32 nm, respectively Parreiras and Martins (2015). Crystalline anisotropy was considered negligible. The following Gilbert damping constants were used for CoFe and NiFe, respectively: $\mathrm{\alpha_{\mathrm{CoFe}}=0.005}$ Weber et al. (2019) and $\mathrm{\alpha_{\mathrm{NiFe}}=0.0095}$ Hrabec et al. (2016). We define the quadratic and biquadratic interlayer exchange interactions using custom expressions for the exchange field and the corresponding energy density. These expressions are constructed to reproduce the exchange energy given in Eq. (1) (the third and fourth terms on the right-hand side) and to ensure that the resulting magnetization dynamics are consistent with the LLG equation. We define the quadratic and biquadratic exchange fields as (see SM):
$\mathrm{\boldsymbol{\mathrm{H}}_{\mathrm{quad}}=-\frac{J_{\mathrm{q}}}{M_{\mathrm{S1}}M_{\mathrm{S2}}}\widetilde{\boldsymbol{\mathrm{M}}}}$, (7a)
$\mathrm{\boldsymbol{\mathrm{H}}_{\mathrm{biquad}}=-\frac{2J_{\mathrm{bq}}}{(M_{\mathrm{S1}}M_{\mathrm{S2}})^{2}}(\widetilde{\boldsymbol{\mathrm{M}}}\cdot\boldsymbol{\mathrm{M}})\widetilde{\boldsymbol{\mathrm{M}}}}$. (7b)
and corresponding energy density terms as:
$\mathrm{\mathscr{E}_{\mathrm{quad}}=-\mathrm{\boldsymbol{\mathrm{M}}}}\cdot\mathrm{\boldsymbol{\mathrm{H}}}_{\mathrm{quad}}.$ (8a)
$\mathrm{\mathscr{E}_{\mathrm{biquad}}=-\frac{1}{2}\mathrm{\boldsymbol{\mathrm{M}}}}\cdot\mathrm{\boldsymbol{\mathrm{H}}}_{\mathrm{biquad}}.$ (8b)
Here, $\widetilde{\boldsymbol{M}}$ is the swapped magnetization defined as $\widetilde{\boldsymbol{M}}(x,y,z,t)=\boldsymbol{M}(x,y,-z,t)$. For the simulations, the saturation magnetizations and exchange coupling constants were taken from macrospin model fits to the STFMR spectra.
We first performed magnetostatic simulations to generate the hysteresis loop. We started the simulation with a static magnetic field of +200+200 mT along the long axis, swept down to 200-200 mT in steps of $2$ mT, and ramped up back to +200+200 mT to complete the hysteresis loop. The built-in relaxation function in MuMax3 was used to find the equilibrium state of the system. The simulated hysteresis loops show good agreement with both the VSM measurements and the macrospin model, as illustrated in Fig. 1(d) for a Ru thickness of $0.8\,\mathrm{nm}$. MuMax3 simulations of the equilibrium angles of the ferromagnetic layers with respect to the external field, $\phi_{\mathrm{a}}$ (CoFe) and $\phi_{\mathrm{b}}$ (NiFe), are plotted in Fig. 1(e) as a function of the applied field during field ramp-down. Note that -180° and 180° correspond to the same magnetization direction. In some regions, the blue data points (magnetization direction of CoFe) are obscured by the overlapping orange data points (magnetization direction of NiFe). This plot provides direct insight into the transitions observed in the magnetostatic measurements and calculations shown in Fig. 1(d). In particular, the transition from configuration ‘1’ to configuration ‘2’ corresponds to a spin-flop transition, whereas the transition from configuration ‘3’ to configuration ‘4’ corresponds to a spin-flip transition.
Figure 4: (a) Coupling gap $g$ (blue) and degeneracy field $H_{\mathrm{g}}$ (green) obtained from macrospin calculations (solid lines) and micromagnetic simulations (dashed lines) as a function of the quadratic exchange constant $J_{\mathrm{q}}$ at a fixed biquadratic exchange constant of $J_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$. The degeneracy field $H_{\mathrm{g}}$ is defined as the magnetic field at which the frequency separation between the high- and low-frequency modes is minimized, yielding the coupling gap $g$. The uncertainty (error bar) in the simulated $H_{g}$ arises from the finite field resolution, given by the field step size (2 mT in this case). The uncertainty in the simulated $g$ is determined by the finite frequency resolution of the FFT, set by the simulation runtime (100 MHz for 10 ns), and is multiplied by $\sqrt{2}$ to account for error propagation. (b) Same as (a), but shown as a function of the biquadratic exchange constant $J_{\mathrm{bq}}$ at a fixed quadratic exchange constant of $J_{\mathrm{q}}=29.8\penalty 10000\ \mathrm{kJ/m^{3}}$. (c–d) The b coefficient in the sublinear power-law dependence on $J_{\mathrm{bq}}$ and $J_{\mathrm{q}}$, respectively.
Dynamic simulations using MuMax3 reveal the phase relations between the ferromagnetic layers in the dynamic spectra. Figures 2(c,d) show simulated spectra for Ru thicknesses of 0.8 nm and 0.4 nm, corresponding to the experimental STFMR measurements in Figs. 2(a,b). We started the simulation with a static magnetic field of +200 mT and swept down to 0 mT in steps of 1 mT, with the initial magnetization state at each field taken from magnetostatic simulations. A sinc pulse with an amplitude of $0.1\penalty 10000\ \mathrm{mT}$ and a cutoff frequency of 25 GHz was applied along all three spatial axes to simultaneously excite both the acoustic and optical modes Subedi et al. (2025), while a static external field was applied along the long axis of the sample. The RF field direction and phase were chosen to minimize numerical artifacts. At each field step, the dynamics were evolved for 10 ns, and the total magnetization of the coupled layers was Fourier transformed to obtain the spectra. Macrospin model calculations are overlaid on the simulations and show excellent agreement. The spectra show two hybridized magnon branches similar to the experimentally obtained spectra, one branch appearing more prominent than the other, consistent with the general observation that the acoustic mode is readily excited whereas the optical mode is more difficult to excite, particularly in the saturated regime Xiao et al. (2022). Importantly, the RF field direction and phase do not affect the spectra (see SM).
We further analyzed the FFT phase to determine the magnon mode relations at different fields. At 4 mT [configuration ‘3’ in Fig. 1(d,e); see Fig. 3(b)], when the layers are antiferromagnetically aligned, the FFT magnitude (top panel) exhibits two peaks at 1.16 GHz and 7.68 GHz in both ferromagnetic layers, marked by vertical lines. At the lower frequency branch (1.16 GHz), the CoFe and NiFe phases (bottom panel) are $151.6^{\circ}$ and $151.3^{\circ}$, marked by solid horizontal lines, differing by only $0.3^{\circ}$, consistent with an in-phase acoustic mode. At the high-frequency branch (7.68 GHz), the phases are $14.4^{\circ}$ (CoFe) and $157.3^{\circ}$ (NiFe), marked by dashed horizontal lines, differing by $171.7^{\circ}$, characteristic of an out-of-phase optical mode.
At the transition field $H_{\mathrm{0,afm}}=12\penalty 10000\ \mathrm{mT}$ [transition point from configuration ‘2’ to configuration ‘3’ in Fig. 1(d,e); Fig. 3(c)], the phase differences remain close to $1^{\circ}$ (acoustic) and $189^{\circ}$ (optical). Near the degeneracy point, the FFT phase exhibits a rapid variation. For this reason, we performed simulations with higher magnetic-field resolution, using a step size of 0.2 mT. At 37.4 mT [Fig. 3(d)], we find an avoided level crossing and a gap in the spectrum, where the phase differences increase to $40^{\circ}$ and $230^{\circ}$, respectively, neither close to what we would expect from an acoustic or optical mode. The modes are no longer purely acoustic or optical in nature, indicating hybridization of the modes. At the second transition field $H_{\mathrm{0,ff}}=52\penalty 10000\ \mathrm{mT}$ [transition point from configuration ‘1’ to configuration ‘2’ in Fig. 1(d,e); Fig. 3(e)], the phase differences switch to $188^{\circ}$ (low frequency) and $6^{\circ}$ (high frequency), corresponding to an interchange of acoustic and optical character of the two modes. In the saturated regime [configuration ‘1’; Fig. 3(f)], at 100 mT the phase differences are $197^{\circ}$ and $7^{\circ}$, again indicating optical and acoustic modes, respectively. Finally, spatially resolved simulations at 12 mT [Fig. 3(g)] show a uniform phase within each layer, with no internal phase variation.
The opening of a gap at the degeneracy field is the signature of a strong coupling between the acoustic and optical modes. The results obtained by simulations support the experimental results (Fig. 2), corroborating a strong coupling between the modes. We argue that this strong coupling between acoustic and optical magnons is facilitated by exchange interaction induced by broken intrinsic symmetry in the studied synthetic ferrimagnet L D Landau and Lifshitz (1977).
The macrospin model and micromagnetic simulations further reveal the dependence of the coupling gap $g$ and the corresponding field $H_{\mathrm{g}}$ on the quadratic and biquadratic exchange interactions. As shown in Figs. 4(a,b), $H_{\mathrm{g}}$ increases linearly with both $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$. In contrast, the gap size $g$ exhibits a sublinear power-law dependence on $J_{\mathrm{q}}$, with the scaling behavior governed by the strength of $J_{\mathrm{bq}}$.
The dependence of $g$ on $J_{\mathrm{q}}$ can be modeled as:
$g=a\,J_{\mathrm{q}}^{\,b}+c,$ (9)
where the fitting parameters $a$, $b$, and $c$ depend on the biquadratic exchange interaction. For $J_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$, we obtain a fitted exponent of $b=0.79$ [Fig. 4(a)]. More generally, the exponent $b$ increases with increasing $J_{\mathrm{bq}}$, starting from $b=0.5$ in the absence of biquadratic exchange ($J_{\mathrm{bq}}=0$) and approaching a linear scaling regime ($b=1$) at large $J_{\mathrm{bq}}$, as shown in Fig. 4(c).
The dependence of the gap $g$ on the biquadratic exchange interaction $J_{\mathrm{bq}}$ is regime dependent. For $J_{\mathrm{q}}<2J_{\mathrm{bq}}$, where a fully antiferromagnetic alignment is not energetically accessible, $g$ follows a sublinear power-law dependence on $J_{\mathrm{bq}}$, with the scaling exponent determined by $J_{\mathrm{q}}$. In contrast, for $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, where a fully antiferromagnetic state becomes accessible [see Eq. (2b)], the gap $g$ decreases with increasing $J_{\mathrm{bq}}$ and deviates from the sublinear power-law behavior.
For $J_{\mathrm{q}}=30\penalty 10000\ \mathrm{kJ/m^{3}}$, we obtain an approximately linear dependence of $g$ on $J_{\mathrm{bq}}$ in the regime $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, with a fitted exponent of $b=1.02$ [Fig. 4(b)]. More generally, the exponent $b$ increases with increasing $J_{\mathrm{q}}$, starting from $b=0.5$ in the absence of quadratic exchange ($J_{\mathrm{q}}=0$) and approaching a linear scaling regime ($b=1$) at large $J_{\mathrm{q}}$, as summarized in Fig. 4(d). Further details are provided in the SM.
IV Summary
In summary, we demonstrated strong coupling between acoustic and optical modes in a synthetic ferrimagnet, CoFe(10 nm)/Ru($x$ nm)/NiFe(10 nm), stabilized by broken intrinsic symmetry in an in-plane geometry. Macrospin model and micromagnetic simulations reveal that both the gap field and gap size depend sensitively on the quadratic and biquadratic exchange interactions, $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$. While $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ can in principle be extracted from magnetostatics via Eqs. (3a) and (3b), this approach lacks precision and fails for $J_{\mathrm{q}}<2J_{\mathrm{bq}}$ where $H_{0,\mathrm{afm}}$ does not exist. By contrast, the macrospin model developed here provides a robust route to parameter extraction by fitting directly to spin-dynamics spectra. Moreover, micromagnetic simulations enable clear distinction between acoustic and optical modes and, through phase analysis, reveal hybridization at the degeneracy field. Our micromagnetic framework, incorporating both quadratic and biquadratic interactions, is broadly applicable to unconventional geometries, including artificial spin-ice structures Sultana et al. (2025); Dion et al. (2024), where the macrospin models break down. Furthermore, our study introduces intrinsic symmetry-breaking as a mechanism to tune magnon-magnon coupling in an in-plane geometry, which offers promising prospects for the development of reconfigurable magnonic devices, such as tunable filters, non-reciprocal components, and coherent magnon-based logic elements. The ability to engineer and control hybridized modes through intrinsic symmetry-breaking can significantly enhance the design flexibility of future on-chip spintronic and magnonic circuits.
Acknowledgments
This research was supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The authors acknowledge the use of facilities and instrumentation supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824.
The data supporting this study’s findings are at [?].